# Instance Segmentation of Scene Sketches Using Natural Image Priors

### v1

@@ Appendix C Additional Qualitative Comparisons
Figure 21: Qualitative comparison of segmentation on filtered SketchyScene dataset. We prompt the semantic segmentation model with ground truth class labels; however, it struggles to locate the objects due to the significant deviation of this sketch style from the data it was trained on. Our performance on both complete and filtered scenes are equally as robust, detecting and segmenting object instances precisely.
Input OpenVocab Ours
@@
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
@@
Figure 22: Qualitative comparison of segmentation on filtered SketchAgent dataset. We prompt the semantic segmentation model with ground truth class labels; however, OpenVocab struggles to accurately identify instances of the correct class, often assigning multiple class labels to the same object, as indicated by the color gradients. Sometimes it is unable to label any sketch pixels in the scene (row 5). In contrast, our method effectively segments object instances, ensuring clear separation and consistent labeling.
Input OpenVocab Ours
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
@@
Figure 23: Qualitative comparison of segmentation on filtered CLIPasso dataset. We prompt the semantic segmentation model with ground truth class labels; however, OpenVocab struggles to accurately identify instances of the correct class, often assigning multiple class labels to the same object, as indicated by the color gradients. Sometimes it is unable to label any sketch pixels in the scene (row 4). In contrast, our method effectively segments object instances, ensuring clear separation and consistent labeling.
Figure 24: ours on SketchyScene inputs (first row), clipasso two styles (second row), sketchagent(last row)

### v2

@@ Appendix G Additional Qualitative Comparisons
Figure 24. Example sketches from our InstantStyle dataset. These sketches are derived from Visual Genome  (2017) containing 5 to 10 annotated objects. For visual clarity, we mask unsegmented regions in the generated sketches. This dataset contains 53 new categories that are not part of SketchyScene’s 45 classes, and contain 1068 sketches in total.
Input SketchyScene Grounded SAM Automatic SAM Ours
Figure 25. Qualitative comparison of instance segmentation methods on the SketchyScene dataset. Our method surpasses SketchyScene, Grounded SAM and Automatic SAM baselines by delivering fine-grained, semantically consistent segmentations with precise boundaries. In the urban scene (row 1), our approach accurately segments all object instances, cleanly separating the sun, tree, building, and character from each other. For the park scenes (rows 2 and 3), it captures intricate details, such as the dog’s face and picnic items, which are either missed or over-segmented by other methods. In the residential and cottage scenes (rows 4 and 5), our method effectively delineates repetitive patterns like fences and handles dense objects like trees, preserving structural integrity where other approaches struggle. These results highlight the robustness of our method in managing complex and detailed sketches.
Input SketchyScene Grounded SAM Automatic SAM Ours
Figure 26. Qualitative comparison of instance segmentation methods on the SketchAgent dataset. Both SketchyScene, Grounded SAM and Automatic SAM struggle to segment these abstract sketches, which differ significantly from their training data of clipart-like and real objects. Our method successfully segments individual instances while maintaining object boundaries, even in challenging cases like the last row where objects overlap with a grid-patterned picnic blanket.
Input SketchyScene Grounded SAM Automatic SAM Ours
Figure 27. Qualitative comparison of instance segmentation methods on the CLIPasso base dataset. Our method successfully detects both large and small objects with ambiguous openings in their silhouettes, whereas baseline methods either merge multiple instances into one or fail to detect them entirely.
Input SketchyScene Grounded SAM Automatic SAM Ours
Figure 28. Qualitative comparison of instance segmentation methods on the InstantStyle dataset. This dataset poses a significantly greater challenge compared to the others due to its increased complexity, including diverse perspectives, intricate textures, and frequent occlusions. Despite these difficulties, our method effectively locates object instances, such as the glass of water in row 2, the fork in row 3, and the food in dishes and bottles in row 4. Even with ambiguous shapes, as shown in row 5, our method outperforms GroundedSAM by successfully segmenting the umbrella on the right separately from the person holding it. In the final row, our approach demonstrates its capability by accurately segmenting both the pile of clothes on the couch and the couch itself.
Input Bourouis et al. SketchSeger Ours
Refer to caption Refer to caption Refer to caption Refer to caption
cat, chair, dog, people, sofa, table
Refer to caption Refer to caption Refer to caption Refer to caption
cat, chair, dog, people, table
Refer to caption Refer to caption Refer to caption Refer to caption
bench, cloud, dog, flower, grass, people, sun, tree
Refer to caption Refer to caption Refer to caption Refer to caption
car, chicken, cloud, cow, fence, grass, house, sun, tree
Refer to caption Refer to caption Refer to caption Refer to caption
bench, duck, moon, star, tree, others
@@
Figure 29. Qualitative comparison of segmentation on filtered SketchyScene dataset. We prompt Bourouis et al. ’s method with ground truth labels and use confidence threshold of 0.01 to ensure all sketch pixels are segmented, while using SketchSeger as-is since it does not accept input labels. However, Bourouis et al. struggles to accurately identify instances of the correct class, often assigning multiple class labels to the same object, as indicated by the color gradients. SketchSeger also fails to provide clean segmentations for many scenes. In contrast, our method effectively segments object instances, ensuring clear separation and consistent labeling.
Input Bourouis et al. SketchSeger Ours
Refer to caption Refer to caption Refer to caption Refer to caption
basket, bucket, cat, cloud, grass, sun
Refer to caption Refer to caption Refer to caption Refer to caption
cloud, dog, duck, grass, house, people, sun, tree
Refer to caption Refer to caption Refer to caption Refer to caption
car, cloud, fence, house, people, sun, tree
Refer to caption Refer to caption Refer to caption Refer to caption
flower, people, tree
Refer to caption Refer to caption Refer to caption Refer to caption
bench, cat, dog, flower, grass, people, picnic rug, tree
@@
Figure 30. Qualitative comparison of segmentation on filtered SketchAgent dataset. We prompt Bourouis et al. ’s method with ground truth labels and use confidence threshold of 0.01 to ensure all sketch pixels are segmented, while using SketchSeger as-is since it does not accept input labels. However, Bourouis et al. struggles to accurately identify instances of the correct class, often assigning multiple class labels to the same object, as indicated by the color gradients. SketchSeger also fails to provide clean segmentations for many scenes. In contrast, our method effectively segments object instances, ensuring clear separation and consistent labeling.
Input Bourouis et al. SketchSeger Ours
Refer to caption Refer to caption Refer to caption Refer to caption
cloud, grass, mountain, people, sun, tree
Refer to caption Refer to caption Refer to caption Refer to caption
cat, chicken, cloud, cow, dog, duck, fence, grass, horse, house, pig, rabbit, sheep, tree
Refer to caption Refer to caption Refer to caption Refer to caption
chicken, cloud, cow, duck, grass, horse, house, pig, sheep, tree, truck
Refer to caption Refer to caption Refer to caption Refer to caption
apple, basket, butterfly, flower, grass, picnic rug, pig, sun, tree
Refer to caption Refer to caption Refer to caption Refer to caption
dog, duck, grass, moon, star, tree
Figure 31. Qualitative comparison of segmentation on filtered CLIPasso dataset. We prompt Bourouis et al. ’s method with ground truth labels and use confidence threshold of 0.01 to ensure all sketch pixels are segmented, while using SketchSeger as-is since it does not accept input labels. However, Bourouis et al. struggles to accurately identify instances of the correct class, often assigning multiple class labels to the same object, as indicated by the color gradients. SketchSeger also fails to provide clean segmentations for many scenes. In contrast, our method effectively segments object instances, ensuring clear separation and consistent labeling.
Input Bourouis et al. SketchSeger Ours
Refer to caption Refer to caption Refer to caption Refer to caption
person, suitcase, train, umbrella
Refer to caption Refer to caption Refer to caption Refer to caption
cloud, horse, mountain, river, tree
Refer to caption Refer to caption Refer to caption Refer to caption
bread, chair, cup, knife, pizza, table
Refer to caption Refer to caption Refer to caption Refer to caption
book, candle, cell phone, table
Refer to caption Refer to caption Refer to caption Refer to caption
bed, clock, couch, person, pillow
Refer to caption Refer to caption Refer to caption Refer to caption
bed, book, clock, mug, person, pillow, table, vase
Refer to caption Refer to caption Refer to caption Refer to caption
bench, bird, bush, cat, cloud, hat, person, tree
Figure 32. Qualitative comparison of segmentation on filtered Zhang et al. dataset. We prompt Bourouis et al. ’s method with ground truth labels and use confidence threshold of 0.01 to ensure all sketch pixels are segmented, while using SketchSeger as-is since it does not accept input labels. Our method accurately segments object instances across diverse scenes, including city life (first and last rows), natural environments (second row), and indoor settings (third to sixth rows). In contrast, baseline methods often struggle to produce clean segmentations for individual objects.
Input Bourouis et al. SketchSeger Ours
Refer to caption Refer to caption Refer to caption Refer to caption
person, dining table, cake
Refer to caption Refer to caption Refer to caption Refer to caption
person, bottle, laptop
Refer to caption Refer to caption Refer to caption Refer to caption
bowl, chair, cup, dining table, sandwich
Refer to caption Refer to caption Refer to caption Refer to caption
book, dining table, cake, person
Refer to caption Refer to caption Refer to caption Refer to caption
cake, cup, dining table, person, spoon
Figure 33. Qualitative comparison of segmentation on filtered InstantStyle dataset. We prompt Bourouis et al. ’s method with ground truth labels and use confidence threshold of 0.01 to ensure all sketch pixels are segmented, while using SketchSeger as-is since it does not accept input labels. Compared to Bourouis et al., which sometimes fails to delineate smaller or overlapping objects, and SketchSeger, which often merges semantically distinct regions or misses foreground elements, our method produces cleaner, more precise instance boundaries and better preserves object semantics—particularly in scenes with clutter, occlusion, or fine-grained details.